OpenStudy (anonymous):

write the product in standard form. (2+3i)(2-i)

myininaya (myininaya):

To multiply (a+b)(c+d) you use distributive property a couple of times A(c+d)+b(c+d)

myininaya (myininaya):

That A is suppose to be a

myininaya (myininaya):

a(c+d)+b(c+d) *

OpenStudy (lgbasallote):

^os bug

myininaya (myininaya):

ac+ad+bc+bd

myininaya (myininaya):

Also recall \[i \cdot i= i^2=-1\]